A bullet of mass m is fired at speed v0 into a wooden block of mass M. The bullet instantaneously comes to rest in the block. The block with the embedded bullet slides along a horizontal surface with a coefficient of kinetic friction m.
1
Expert's answer
2020-09-08T09:12:26-0400

Let us determine the momentum of the bullet. It is "mv_0." Next, we should determine the velocity "v" of the system "bullet+block". According to the momentum conservation law, we get "mv_0 = (M+m)v", so "v = \dfrac{m}{m+M}v_0" . Next, we should calculate the kinetic energy of the system "bullet+block" at the beginning of their simultaneous motion. The energy will be

"E_0 = \dfrac{(m+M)v^2}{2} = \dfrac{m^2v_0^2}{2(m+M)}."

The system will stop when the modulus of the friction force work will be equal to "E_0," so "E_0=|A_{fr}|" , where "|A_{fr}| = \mu NS = \mu (m+M)gS." Therefore, the distance, after that the system will stop, is "S = \dfrac{E_0}{\mu (m+M)g} = \dfrac{m^2v_0^2}{2\mu (m+M)^2g}" .